Lane works in Quebec and is paid on a semi-monthly basis. He has the following types...

  1. Lane works in Quebec and is paid on a semi-monthly basis. He has the following types of earnings & benefits. Determine which are pensionable and insurable and then calculate the CPP and EI deductions.

  1. Statutory Holiday Pay $120
  2. Regular $1200
  3. Employer paid Parking $30
Answer #1

Calculation of CPP

Total pensionable income = $1200 + $120 + $ 30 = $ 1350. (Step 1)

Basic pay period exemption for 2020 (Semi- mostly basis) =   $145.83 (Step 2)

Step 1 - Step 2 = $1350 - $ 145.83 = $ 1204.17 (Step 3)   

Step 3 * CPP contribution rate 2020. = $1204.17 * 5.25 %. =. $ 63.2189

So CPP for the period = $ 63.2189

Calculation of EL

Insurable Amount = $ 1200 + $ 120 = $ 1320 (Step 1)

Employee EL premium rate for 2020 = 1.58 % (Step 2)

EL for the period = 1320 * 1.58% = $ 20.856